OpenStudy (anonymous):

convert the polar equation to rectangular coordinates: tan(theta)=8

OpenStudy (anonymous):

is the answer y=8x??

OpenStudy (turingtest):

no, because it seems to be a circle

OpenStudy (anonymous):

x=r*cos(Θ) y=r*sin(Θ) x^2 + y^2 = r^2

OpenStudy (turingtest):

and tan(theta)=y/x

OpenStudy (turingtest):

I guess it is y=8x then?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

ok thx for replying